OpenStudy (anonymous):

three boys ride in 3meter seesaw. two of the smaller boys ride on one side, and the third bigger boy on the end of the other side. what is the mass of the bigger boy, if the 35kg boys sits on the end of the seesaw while the second 30 kg boy sits 1.0 meter away from the fulcrum.

OpenStudy (mathstudent55):

This is a Physics problem on conservation of momentum.

OpenStudy (anonymous):

can you help me?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

x(1.5 m) = 35(1.5 m) + 30(1.0 m) Solve for x

OpenStudy (anonymous):

okie.. thanks :)

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

5.5?

OpenStudy (mathstudent55):

55 kg, isn't it?

OpenStudy (mathstudent55):

1.5x = 52.5 + 30 1.5x = 82.5 x = 55

OpenStudy (anonymous):

:)